# New Characteristics and Modeling of 6G Channels: A Unified Channel Model towards Standardization

### v1

@@ III-D Simulation of the unified channel model
Figure 7: CIR of new frequency communication under unified channel model.
Fig. 4 illustrates the RCS modeling characteristics of the sensing targets in the simulation platform. The simulation includes a UAV with dimensions of $347\times 283\times 107$ mm, which is observed in a monostatic sensing setup. As previously discussed, the RCS of the UAV can be assumed isotropic in the horizontal plane and is modeled using a Gaussian distribution $N(\mu,\sigma^{2})$. The values of $\mu$ and $\sigma$ are referenced from the report of the TSG RAN#118 meeting [16].
@@
Fig. 6 illustrates the impact of RIS configuration on the receiving SNR of the concatenated Tx-RIS-Rx channel in the unified channel model framework. The positions of the transmitter, receiver, and RIS are fixed, while the RIS codebook configuration is varied, including specular reflection (codebook set to all zero), 1-bit phase shift, and continuous phase shift configurations. The side length of each RIS element is set to $\lambda/2$, where $\lambda$ is the wavelength of the corresponding frequency band. The simulation result shows that the receiving SNR increases with the number of RIS elements, and continuous phase shift provides better performance compared to the 1-bit phase shift and specular reflection configurations. This indicates that the proposed simulation framework effectively captures the influence of RIS on channel performance.

### v2

@@ III-D Simulation of the unified channel model
Figure 7: Comparison of Gini coefficients between 3GPP model and unified model in new frequency band.
Fig. 4 illustrates the RCS modeling characteristics of the sensing targets in the simulation platform. The simulation includes a UAV with dimensions of $347\times 283\times 107$ mm, which is observed in a monostatic sensing setup. As previously discussed, the RCS of the small UAV can be assumed isotropic in the horizontal plane and is modeled using a Gaussian distribution $N(\mu,\sigma^{2})$. The values of $\mu$ and $\sigma$ are referenced from the report of the TSG RAN#118 meeting [15].
@@
Fig. 6 illustrates the impact of RIS configuration on the received SNR of the concatenated Tx-RIS-Rx channel within the unified channel model framework. The positions of the transmitter, receiver, and RIS are fixed, while the RIS codebook configuration is varied, including specular reflection (codebook set to all zeros), 1-bit phase shift, and continuous phase shift configurations. The simulation results show that for 1-bit and continuous phase shift configurations, the received SNR increases with the number of RIS elements, and continuous phase shift achieves better performance than 1-bit phase shift. For the specular reflection configuration, since no phase compensation is applied, the signals reflected by different RIS elements experience phase mismatches. As the RIS array size increases, the coherent addition of signals leads to fluctuations in the received power. This indicates that the proposed simulation framework effectively captures the influence of RIS on channel performance.